OpenStudy (anonymous):

what is 2 3/4 divided by -1/8

OpenStudy (anonymous):

please give specific explanations please

OpenStudy (anonymous):

\[\frac{2\frac{3}{4}}{-\frac{1}{8}}\]\[\frac{11}{4} \times -\frac{8}{1}\]\[\frac{-88}{4}\]\[-22\]

OpenStudy (anonymous):

thanks you are very helpful